Two elecctrical charges of 500 m(permeability)C each exert a force of 3.2 N on each other. How far apart are they? Next problem. Two electrical charges placed 100 m apart exert a Force of 6200 N on each other. What is the charge? Can anyone give the answer to these physics questions I’m stuck on?
1
Expert's answer
2020-04-10T08:52:30-0400

Coulomb's law:

"F_C=k_e \cdot \frac{q_1 \cdot q_2}{r^2}",

where "F_C" - Coulomb force, N; "q_1, \; q_2" - electrical charges, C; "r" - distance, m; "k_e=9 \cdot 10^9 \; \frac{N \cdot m^2}{C^2}" - Coulomb's constant. If electrical charges are equal ("q_1=q_2=q"):

"F_C=k_e \cdot \frac{q^2}{r^2}"

1) "r=q \cdot \sqrt{\frac{k_e}{F_C}}"

"r=500 \cdot 10^{-3} \cdot \sqrt{\frac{9 \cdot 10^9}{3.2}} \approx 26516 \; m \approx 26.5 \; km"

2) "q=r \cdot \sqrt{\frac{F_C}{k_e}}"

"q=100 \cdot \sqrt{\frac{6200}{9 \cdot 10^9}} \approx 83 \cdot 10^{-3} \; C = 83 \; mC"